# Quintessential early dark energy

### v2

@@ 6 Growth of structures
$\mathcal{Q}=\frac{\mathcal{B}(t,k,k^{\prime})}{\mathcal{P}(t,k)\mathcal{P}(t,k^{\prime})+\mathcal{P}(t,k^{\prime})\mathcal{P}(t,k^{\prime\prime})+\dots}\,,$ (6.17)
because it is scale and time independent to lowest order in nonlinear perturbation theory. Fig. 6 shows the nature of the reduced bispectrum at $z=0$ for the triangle configuration given by $k/k^{\prime}=1$ with $k=0.01h{\rm Mpc}^{-1}$ (left figure) and $k/k^{\prime}=0.2$ with $k=0.01h{\rm Mpc}^{-1}$ (right figure) as a function of the angle $\theta$ between the wave vectors $\vec{k}$ and $\vec{k}^{\prime}$ given by $\hat{k}.\hat{k}^{\prime}=\cos\theta$. We don’t see any difference int the reduced bispectrum for the MSE potential and $\Lambda$CDM model.
@@ 7.1.1 CMB
$R(z_{\star})$ $=$ $(1+z_{\star})D_{A}(z_{\star})\frac{{\sqrt{\Omega_{\mathrm{m}0}H_{0}^{2}}}}{{c}}$ (7.1)
$=$ $\sqrt{\Omega_{\mathrm{m}0}}\int_{0}^{z_{\star}}\frac{\mathrm{d}z}{E(z)}\,$ (7.2)
$l_{\rm A}$ $=$ $\frac{\pi}{\theta_{\star}}=\frac{{\pi(1+z_{\star})D_{A}(z_{\star})}}{{r_{s}(z_{\star})}}\ ,$ (7.3)
where, $z_{\star}$ is the decoupling redshift, $r_{s}$ signifies the sound horizon as defined in Eq. (5.2) and $D_{A}$ corresponds to the angular diameter distance defined in Eq. (5.3). The baryon-photon ratio $R_{\rm b}$, defined in Eq. (5.5), can be related to CMB temperature $T_{\rm CMB}=2.7255K$ as
$R_{\rm b}=31500\omega_{\rm b}\left(\frac{T_{\rm CMB}}{2.7K}\right)^{-4}\,.$ (7.4)
@@
$\Omega_{\mathrm{r}0}=\frac{\Omega_{\mathrm{m}0}}{1+z_{\rm eq}}\,,$ (7.5)
@@
$z_{\rm eq}=25000\omega_{\mathrm{m}}\left(\frac{T_{\rm CMB}}{2.7K}\right)^{-4}\,,$ (7.6)
@@ 7.2 Data combinations
For a detail analysis and comparison with other scenario we perform the data analysis for the SE (3.1), MSE (4.1), axionlike (5.12) and power law (5.13) potentials along with the standard $\Lambda$CDM model. The different data combinations are CMB+BAOi+SNIaj+Hubble, where subscript BAOi stands for BAO and DBAO while SNIaj stands for Pantheon$+$, Union3 and D5. So, we have six combinations of different observational data. We call CMB+DBAO+Pantheon$+$
@@ 7.3.2 SE potential
Apart from the parameters mentioned in Tab. 2 for $\Lambda$CDM the model parameters for the SE potential (3.1) are $\{\mu,\;V_{0}\}$. We take $\log_{10}V_{0}$ as the model parameter instead of $V_{0}$. $V_{0}$ sets the scale at which EDE starts decaying after reaching its maximum value. In the literature this scale is generally represented as the redshift $z_{\rm c}$. $\mu$ determines the peak value of EDE at $z_{\rm c}$. $\log_{10}V_{0}$ also has some effect on this. The fractional energy density at the peak value of EDE at $z_{\rm c}$ is denoted by $f_{\rm EDE}$ in the literature. In our analysis instead of using the standard $f_{\rm EDE}$ and $z_{\rm c}$ parameters we use the parameters $\mu$ and $\log_{10}V_{0}$. Also, we do not use any parametrisation of the scalar field energy density as suggested in Ref. [41] rather we completely rely on solving the system numerically. The priors for the common parameters with the $\Lambda$CDM model are same as Tab. 2. The priors for the parameters $\Big{\{}\mu,\;\log_{10}V_{0}\Big{\}}$ are $\Big{\{}[0.1,5.0],\;[-2.0,18.0]\Big{\}}$.
@@ 7.3.3 MSE potential
For the MSE potential (4.1) the extra parameters, other than the ones we have for $\Lambda$CDM model, would be $\phi_{0}$ and $V_{0}$. Here also we take $\log_{10}V_{0}$ as the model parameter instead of $V_{0}$. The prior on the parameters $\Big{\{}\phi_{0},\;\log_{10}V_{0}\Big{\}}$ are $\Big{\{}[0.25,5],\;[-2,18]\Big{\}}$ for $n=6$ and $\Big{\{}[0.5,5],\;[-2,18]\Big{\}}$ for $n=10$.
@@ 7.3.4 Axionlike potential
Similar to the SE and MSE potential for the axionlike potential (5.12) also we have two extra parameter compared to the $\Lambda$CDM model. The extra parameters are $\Big{\{}f,\;\log_{10}V_{0}\Big{\}}$ and their priors are $\Big{\{}[0.01,0.3],\;[-2,18]\Big{\}}$.
@@ 7.3.5 Power law potential
For the power potential (5.13) also we have two more parameters than the $\Lambda$CDM model and they are $\Big{\{}f,\;\log_{10}V_{0}\Big{\}}$. The priors of these two parameters are $\Big{\{}[0.2,3];[-5,18]\Big{\}}$.
Table 3: Constraints on the parameters for SE (3.1), MSE (4.1), axionlike (5.12) and power law (5.13) potentials. The initial conditions to solve the Eq. (2.4) are $\{\phi_{i},\phi^{\prime}_{i}\}\equiv\{1.13,0.15\},\{0.9\phi_{0},10^{-10}\},\{\pi f-0.01f,10^{-15}\}\;{\rm and}\;\{f,10^{-15}\}$ for the SE, MSE, axionlike and power law potential respectively.
@@ 7.4 Results
$\Delta\rm AIC$ $=$ $\rm AIC_{\rm EDE}-AIC_{\rm\Lambda CDM}$ (7.10)
$\Delta\rm BIC$ $=$ $\rm BIC_{\rm EDE}-BIC_{\rm\Lambda CDM}\,.$ (7.11)
$\Delta$AIC and $\Delta$BIC tell us about the preference of the model by the observational data in comparison to a reference model. In our case, we have considered $\Lambda$CDM model as the reference model which also has less number of free parameters than the EDE models. From Tab. LABEL:tab:AIC we can infer from the values of $\chi^{2}_{\rm min}$ and $\chi^{2}_{\rm red}$ that the data prefers the $\Lambda$CDM model more than the EDE models as the values of $\chi^{2}_{\rm min}$ and $\chi^{2}_{\rm red}$ are minimum for $\Lambda$CDM except for the data combinations with DESY5. In this regard, it should be noted that even though we have some constraint on the parameters $\{\mu,\;\phi_{0},\;f\}$ the constraints are not very stringent and one has to take care of it while calculating AIC and BIC. In Tab. LABEL:tab:AIC the values of AIC and BIC are listed. EDE models also have similar values of $\chi^{2}_{\rm min}$ and $\chi^{2}_{\rm red}$, except for the data combinations with DESY5, but they have more free parameters. So, for our case it is safe to choose the standard $\Lambda$CDM model as the reference model. Now, larger the values of $\Delta$AIC and $\Delta$BIC lesser is the probability that the data will prefer our model over the standard $\Lambda$CDM model. More precisely, if $20<\Delta\rm AIC,\Delta\rm BIC<$ 2 then the observational data can not distinguish between the models. If $\Delta\rm AIC,\Delta\rm BIC>10$ then the deviation of the considered model from the reference model is decisive. In between these limits the deviation is significant [97]. For the data combinations with Pantheon$+$ the data prefer MSE potential (4.1) with $n=10$ compared to the other EDE potentials. But, for the data combinations with Union3 the data prefer power law potential (5.13) with $n=6$ over the other EDE potentials. Even though there are slight preferences on one EDE potential over other by some data combinations the preference is not very significant. For the data combinations with DESY5 the data show preference towards EDE potentials specially the SE potential (5.13) with $n=10$ and power law potential (5.13) with $n=6$. For DESY5 only MSE potential behaves closer to the $\Lambda$CDM model.
@@
The constraints on the parameters for the SE potential (3.1), shown in Tab. LABEL:tab:constraint, are for $n=6$ and 10. It has been checked that the larger values of $n$ can give rise to larger values in the parameter $h$ (see Tab. 1) and that is why we have chosen larger values of $n$. Smaller values of $n$ can lead to scaling behaviour in the scalar field dynamics just after the redshift $z_{\rm c}$ where $\rho_{\phi}$ scales the background energy density [58]. This behaviour forces the maximum value of EDE density at $z_{\rm c}$ to be very small so that $\rho_{\phi}$ remains small when it scales the background. But we need the EDE energy density to decay very fast just after $z_{\rm c}$ so that it does not affect the late time dynamics of the universe. This can easily be achieved for larger values of $n$ as depicted in the upper figure of Fig. 1. Similar to the SE potential we have chosen larger values $n$ for MSE potential also. One should note that the MSE potential leads to tracker behaviour in the scalar field dynamics during the late time. This tracker behaviour can lead to even less amount of EDE at the redshift $z_{\rm c}$ which may lead to lesser values in $h$ compared to SE potential. Tab. LABEL:tab:constraint and 5 show these behaviour in the values of $f_{\rm EDE}$ and $h$. In fact, from Tab. 5, we see that, for MSE potential, the observationally allowed value of $f_{\rm EDE}$ is almost zero even though theoretically non-zero EDE is possible in MSE potential (see Fig. 4 and Tab. 1). For SE potential the allowed value of $f_{\rm EDE}$ are non zero but small which gives very slight improvement in $h$ compared to MSE potential. But the improvement is not much and the constrained values are similar to the $\Lambda$CDM model. Similar to the SE and MSE potential we also don’t see any improvement in the values of $h$ for the power law potential. On the other hand the axionlike gives rise to higher values in $h$. Here, one should note that, for the oscillatory potentials (axionlike and power law), for the considered data combinations, even though the values of $h$ are larger than that of $\Lambda$CDM model for different data combinations the values are significantly lower than the values obtained in Ref. [40, 41, 42, 43] but very similar to the constraint obtained in [56] for the axionlike potential without considering the SH0ES prior [10]. In our analysis we have not considered the SH0ES prior. Also, we have considered a wide range of the prior for the parameter $\log_{10}V_{0}$ so that we can accommodate the $\Lambda$CDM model within the models with SE, MSE, axionlike and power law potential. We see that the considered observational data has a tendency to go towards $\Lambda$CDM model. The maximum deviation we get for the data combination of DESIUnion. In general the data combination with DESI BAO data prefers larger value of $h$ compared to the other data combinations.
@@ 8 Summary and Conclusion
In this paper we study a quintessential early dark energy model where the EDE and LDE can be achieved through a single scalar field by introducing the MSE potential (4.1). The MSE potential is a modification of the SE potential (3.1) introduced in Ref. [58] in the context of quintessential inflation [62]. For SE potential the late time acceleration is achieved by considering a nonminimal coupling of the scalar field with massive neutrinos [58]. But unlike the SE potential MSE potential can give rise to slow roll in scalar field dynamics during the late time evolution of the universe. In this way we can relate two energy scales with the MSE potential. So, by fixing the energy scale of the scalar field potential at a high energy we can get some desired low energy scale by fixing the parameters suitably (see Eq (4.6)). SE potential leads to scaling behaviour in the scalar field dynamics during late time as shown in Fig. 1. On the other hand the MSE potential lead to tracker behaviour as shown in the top right figure of Fig. 4. Both the potentials have steep nature which can lead to EDE solution because of their very steep nature which also results into rapid decay of EDE energy density (see Fig. 4).
As the SE and MSE potentials successfully produce EDE it is important to constrain the model parameters with cosmological data. In this regard we also consider the axionlike (5.12) and power law (5.13) potentials and compare the constraint values of the model parameters. To do this comparison we first calculated the possible theoretical values of $H_{0}$ by using Eq. (5.10) and following the Ref. [11] which gives a theoretical estimation while using some of the constraint of the $\Lambda$CDM model. This theoretical estimation is given in Tab. 1 and we see that all four potentials can produce larger values of $H_{0}$ than $\Lambda$CDM due to the presence of EDE solutions. We also see that the SE, axionlike and power law potentials can give rise to higher values of $H_{0}$ than the MSE potential. In fact the MSE potential gives lower values of $H_{0}$ than the $\Lambda$CDM model for lower values of $n$. The reason behind this is the existence of tracker behaviour during the late time in the MSE potential which is more prevalent for lower values of $n$. In tracker dynamics the present value of the EoS parameter can be slightly larger than $-1$ which may decrease the value of $H_{0}$. For all other potentials We have added a CC for late time acceleration which makes the dynamics similar to the $\Lambda$CDM from the very recent past.
The tracker behaviour in the MSE potential also plays a crucial role at the perturbation level. We see a suppression in the linear power spectrum for the MSE potential which has been shown in the left figure of Fig. 5. We also get lower value of $\sigma_{8}$ for the MSE potential than the $\Lambda$CDM model for the same initial conditions. So, at the linear perturbation level we can distinguish the MSE potential from the $\Lambda$CDM model. This difference is also very clearly visible in the evolution of $f\sigma_{8}(z)$ (right figure of Fig. 5). It is interesting to note that the deviation from the $\Lambda$CDM model increases as we lower the value of $n$. The reason behind this is the tracker dynamics which becomes more relevant for lower values of $n$. For reduced bispectrum we don’t see any significant deviation from the $\Lambda$CDM model (Fig. 6).
[…]
In summary, in this paper, we introduce a unified model of early and late dark energy. In this regard we have introduces the MSE potential (4.1) which is a modification of the SE (3.1) potential. MSE potential can behave like a CC during late time. During the intermediate stages the scalar field dynamics is a tracker solution. We also compare the model with observation along with the SE, axionlike and power law potential. We see that even though all four of these potential can give rise to EDE solution which, theoretically, may give larger values of $H_{0}$, observationally, with the data sets considered, we don’t see any significant improvement in the values of $H_{0}$ except for axionlike potential where there is a slight improvement. So, even though the MSE potential can relate two different energy scales and unify EDE and LDE the considered observational data do not prefer any finite value of $f_{\rm EDE}$ and there by no improvement in the values of $H_{0}$ compared to the $\Lambda$CDM. In future we shall verify this even further by including other cosmological data and changing the priors of the parameters.
@@
Note Added: While working on the observational data analysis part of this paper we came across the recent paper [98] where the authors also have considered the MSE potential (4.1) but in the context of inflation.

### v3

@@ 6 Growth of structures
$\mathcal{Q}=\frac{\mathcal{B}(t,k,k^{\prime})}{\mathcal{P}(t,k)\mathcal{P}(t,k^{\prime})+\mathcal{P}(t,k^{\prime})\mathcal{P}(t,k^{\prime\prime})+\dots}\,,$ (56)
because it is scale and time independent to lowest order in nonlinear perturbation theory. Fig. 6 shows the nature of the reduced bispectrum at $z=0$ for the triangle configuration given by $k/k^{\prime}=1$ with $k=0.01h{\rm Mpc}^{-1}$ (left figure) and $k/k^{\prime}=0.2$ with $k=0.01h{\rm Mpc}^{-1}$ (right figure) as a function of the angle $\theta$ between the wave vectors $\vec{k}$ and $\vec{k}^{\prime}$ given by $\hat{k}.\hat{k}^{\prime}=\cos\theta$. We don’t see any difference in the reduced bispectrum for the MSE potential and $\Lambda$CDM model.
@@ 7.1.1 CMB
$R(z_{\star})$ $=$ $(1+z_{\star})D_{A}(z_{\star})\frac{{\sqrt{\Omega_{\mathrm{m}0}H_{0}^{2}}}}{{c}}$ (57)
$=$ $\sqrt{\Omega_{\mathrm{m}0}}\int_{0}^{z_{\star}}\frac{\mathrm{d}z}{E(z)}\,$ (58)
$l_{\rm A}$ $=$ $\frac{\pi}{\theta_{\star}}=\frac{{\pi(1+z_{\star})D_{A}(z_{\star})}}{{r_{s}(z_{\star})}}\ ,$ (59)
where, $z_{\star}$ is the decoupling redshift, $r_{s}$ signifies the sound horizon as defined in Eq. (28) and $D_{A}$ corresponds to the angular diameter distance defined in Eq. (29). The baryon-photon ratio $R_{\rm b}$, defined in Eq. (31), can be related to CMB temperature $T_{\rm CMB}=2.7255K$ as
$R_{\rm b}=31500\omega_{\rm b}\left(\frac{T_{\rm CMB}}{2.7K}\right)^{-4}\,.$ (60)
@@
$\Omega_{\mathrm{r}0}=\frac{\Omega_{\mathrm{m}0}}{1+z_{\rm eq}}\,,$ (61)
@@
$z_{\rm eq}=25000\omega_{\mathrm{m}}\left(\frac{T_{\rm CMB}}{2.7K}\right)^{-4}\,,$ (62)
@@ 7.2 Data combinations
For a detail analysis and comparison with other scenario we perform the data analysis for the SE (15), MSE (18), axionlike (38) and power law (39) potentials along with the standard $\Lambda$CDM model. The different data combinations are CMB+BAOi+SNIaj+Hubble, where subscript BAOi stands for BAO and DBAO while SNIaj stands for Pantheon$+$, Union3 and D5. So, we have six combinations of different observational data. We call CMB+DBAO+Pantheon$+$
@@ 7.3.2 SE potential
Apart from the parameters mentioned in Tab. 2 for $\Lambda$CDM the model parameters for the SE potential (15) are $\{\mu,\;V_{0}\}$. We take $\log_{10}V_{0}$ as the model parameter instead of $V_{0}$. $V_{0}$ sets the scale at which EDE starts decaying after reaching its maximum value. In the literature this scale is generally represented as the redshift $z_{\rm c}$. $\mu$ determines the peak value of EDE at $z_{\rm c}$. $\log_{10}V_{0}$ also has some effect on this. The fractional energy density at the peak value of EDE at $z_{\rm c}$ is denoted by $f_{\rm EDE}$ in the literature. In our analysis instead of using the standard $f_{\rm EDE}$ and $z_{\rm c}$ parameters we use the parameters $\mu$ and $\log_{10}V_{0}$. Also, we do not use any parametrisation of the scalar field energy density as suggested in Ref. [41] rather we completely rely on solving the system numerically. The priors for the common parameters with the $\Lambda$CDM model are same as Tab. 2. The priors for the parameters $\Big{\{}\mu,\;\log_{10}V_{0}\Big{\}}$ are $\Big{\{}[0.1,5.0],\;[-2.0,18.0]\Big{\}}$. For $\mu=0$ and $V_{0}\sim\rho_{\rm DE0}$ the SE potential behaves as a CC. Practically, very small values of $\mu$ makes the potential similar to the CC. So, smaller value of $\mu$ and $V_{0}$ can mimic a CC. This has been checked numerically.
@@ 7.3.3 MSE potential
For the MSE potential (18) the extra parameters, other than the ones we have for $\Lambda$CDM model, would be $\phi_{0}$ and $V_{0}$. Here also we take $\log_{10}V_{0}$ as the model parameter instead of $V_{0}$. The prior on the parameters $\Big{\{}\phi_{0},\;\log_{10}V_{0}\Big{\}}$ are $\Big{\{}[0.25,5],\;[-2,18]\Big{\}}$ for $n=6$ and $\Big{\{}[0.5,5],\;[-2,18]\Big{\}}$ for $n=10$. $\phi_{0}$ changes the amount of EDE. We have chosen the prior of $\phi_{0}$ such a way so that $f_{\rm EDE}$ can have maximum value of $\sim 0.5$. For larger values of $n$, e.g., $n=6~{}{\rm or}~{}10$ the MSE potential behaves like a CC during the present time. Also, for very small values of $V_{0}$ the dynamics can be very similar to $\Lambda$CDM.
@@ 7.3.4 Axionlike potential
Similar to the SE and MSE potential for the axionlike potential (38) also we have two extra parameter compared to the $\Lambda$CDM model. The extra parameters are $\Big{\{}f,\;\log_{10}V_{0}\Big{\}}$ and their priors are $\Big{\{}[0.01,0.3],\;[-2,18]\Big{\}}$. The value of $f_{\rm EDE}$ depends on $f$. Similar to the MSE potential for the axionlike potential also we have chosen the prior such a so that $f_{\rm EDE}$ can have maximum value of $\sim 0.5$. Also, similar to the MSE potential axionlike potential also gives a dynamics similar to the $\Lambda$CDM for smaller values of $V_{0}$ and smaller value of $f$.
@@ 7.3.5 Power law potential
For the power potential (39) also we have two more parameters than the $\Lambda$CDM model and they are $\Big{\{}f,\;\log_{10}V_{0}\Big{\}}$. The priors of these two parameters are $\Big{\{}[0.2,3];[-5,18]\Big{\}}$. In this case also $f$ determines the value of $f_{\rm EDE}$ and within this prior range the maximum value of $f_{\rm EDE}$ can be $0.5$. This prior range can give a $\Lambda$CDM like behaviour for smaller value of $V_{0}$ and smaller value of $f$.
Table 3: Constraints on the parameters for SE (15), MSE (18), axionlike (38) and power law (39) potentials. The initial conditions to solve the Eq. (4) are $\{\phi_{i},\phi^{\prime}_{i}\}\equiv\{1.13,0.15\},\{0.9\phi_{0},10^{-10}\},\{\pi f-0.01f,10^{-15}\}\;{\rm and}\;\{f,10^{-15}\}$ for the SE, MSE, axionlike and power law potential respectively.
@@ 7.4 Results
$\Delta\rm AIC$ $=$ $\rm AIC_{\rm EDE}-AIC_{\rm\Lambda CDM}$ (66)
$\Delta\rm BIC$ $=$ $\rm BIC_{\rm EDE}-BIC_{\rm\Lambda CDM}\,.$ (67)
$\Delta$AIC and $\Delta$BIC tell us about the preference of the model by the observational data in comparison to a reference model. In our case, we have considered $\Lambda$CDM model as the reference model which also has less number of free parameters than the EDE models. From Tab. LABEL:tab:AIC we can infer from the values of $\chi^{2}_{\rm min}$ and $\chi^{2}_{\rm red}$ that the data prefers the $\Lambda$CDM model more than the EDE models as the values of $\chi^{2}_{\rm min}$ and $\chi^{2}_{\rm red}$ are minimum for $\Lambda$CDM except for the data combinations with DESY5. In this regard, it should be noted that even though we have some constraint on the parameters $\{\mu,\;\phi_{0},\;f\}$ the constraints are not very stringent and one has to take care of it while calculating AIC and BIC. In Tab. LABEL:tab:AIC the values of AIC and BIC are listed. EDE models also have similar values of $\chi^{2}_{\rm min}$ and $\chi^{2}_{\rm red}$, except for the data combinations with DESY5, but they have more free parameters. So, for our case it is safe to choose the standard $\Lambda$CDM model as the reference model. Now, larger the values of $\Delta$AIC and $\Delta$BIC lesser is the probability that the data will prefer our model over the standard $\Lambda$CDM model. More precisely, if $20<\Delta\rm AIC,\Delta\rm BIC<$ 2 then the observational data can not distinguish between the models. If $\Delta\rm AIC,\Delta\rm BIC>10$ then the deviation of the considered model from the reference model is decisive. In between these limits the deviation is significant [97]. For the data combinations with Pantheon$+$ the data prefer MSE potential (18) with $n=10$ compared to the other EDE potentials. But, for the data combinations with Union3 the data prefer power law potential (39) with $n=6$ over the other EDE potentials. Even though there are slight preferences on one EDE potential over other by some data combinations the preference is not very significant. For the data combinations with DESY5 the data show preference towards EDE potentials specially the SE potential (39) with $n=10$ and power law potential (39) with $n=6$. For DESY5 only MSE potential behaves closer to the $\Lambda$CDM model.
@@
The constraints on the parameters for the SE potential (15), shown in Tab. LABEL:tab:constraint, are for $n=6$ and 10. It has been checked that the larger values of $n$ can give rise to larger values in the parameter $h$ (see Tab. 1) and that is why we have chosen larger values of $n$. Smaller values of $n$ can lead to scaling behaviour in the scalar field dynamics just after the redshift $z_{\rm c}$ where $\rho_{\phi}$ scales the background energy density [58]. This behaviour forces the maximum value of EDE density at $z_{\rm c}$ to be very small so that $\rho_{\phi}$ remains small when it scales the background. But we need the EDE energy density to decay very fast just after $z_{\rm c}$ so that it does not affect the late time dynamics of the universe. This can easily be achieved for larger values of $n$ as depicted in the upper figure of Fig. 1. Similar to the SE potential we have chosen larger values $n$ for MSE potential also. One should note that the MSE potential leads to tracker behaviour in the scalar field dynamics during the late time. This tracker behaviour can lead to even less amount of EDE at the redshift $z_{\rm c}$ which may lead to lesser values in $h$ compared to SE potential. Tab. LABEL:tab:constraint and LABEL:tab:fEDE show these behaviour in the values of $f_{\rm EDE}$ and $h$. In fact, from Tab. LABEL:tab:fEDE, we see that, for MSE potential, the observationally allowed value of $f_{\rm EDE}$ is almost zero even though theoretically non-zero EDE is possible in MSE potential (see Fig. 4 and Tab. 1). For SE potential the allowed value of $f_{\rm EDE}$ are non zero but small which gives very slight improvement in $h$ compared to MSE potential. But the improvement is not much and the constrained values are similar to the $\Lambda$CDM model. Similar to the SE and MSE potential we also don’t see any improvement in the values of $h$ for the power law potential. On the other hand the axionlike gives rise to higher values in $h$. Here, one should note that, for the oscillatory potentials (axionlike and power law), for the considered data combinations, even though the values of $h$ are larger than that of $\Lambda$CDM model for different data combinations the values are significantly lower than the values obtained in Ref. [40, 41, 42, 43] but very similar to the constraint obtained in [56] for the axionlike potential without considering the SH0ES prior [10]. In our analysis we have not considered the SH0ES prior. Also, we have considered a wide range of the prior for the parameter $\log_{10}V_{0}$ so that we can accommodate the $\Lambda$CDM model within the models with SE, MSE, axionlike and power law potential. We see that the considered observational data has a tendency to go towards $\Lambda$CDM model. The maximum deviation we get for the data combination of DESIUnion. In general the data combination with DESI BAO data prefers larger value of $h$ compared to the other data combinations.
@@ 8 Summary and Conclusion
We examine a quintessential EDE model where both EDE and LDE arise from a single scalar field with the MSE potential (18). The MSE potential refines the SE potential (15), initially proposed in [58] in the context of quintessential inflation [62] scenario. While the SE potential follows a scaling solution during the late time (Fig. 1), the MSE potential exhibits tracker dynamics (Fig. 4). Both allows a transient EDE phase but unlike SE potential MSE potential ensures late time acceleration without additional interactions.
We compare MSE with SE (15), axionlike (38), and power law (39) potentials, constraining parameters using cosmological data. Although, theoretically, all four potentials predict a larger $H_{0}$ than $\Lambda$CDM (Tab. 1), the MSE potential, with its tracker behaviour, yields a more conservative increase, and for lower $n$, even predicts $H_{0}$ values below $\Lambda$CDM. This occurs for lower values of $n$ as the scalar field exhibits tracker dynamics even during the recent past making the scalar field EoS larger than $-1$. The MSE potential also affects perturbations, suppressing the linear power spectrum (Fig. 5, left) and lowering $\sigma_{8}$. This suppression is visible in the $f\sigma_{8}(z)$ evolution (Fig. 5, right), particularly for small $n$, though the reduced bispectrum remains close to $\Lambda$CDM (Fig. 6). We see this suppression only for the MSE potential i.e., the scalar field with tracker dynamics.
Observational constraints (Tab. LABEL:tab:constraint) reveal that among the tested EDE potentials, the axionlike model shows the largest deviation in $h$, whereas SE, MSE, and power law potentials remain closer to $\Lambda$CDM. Notably, the DESIUnion data combination exhibits the strongest deviation from $\Lambda$CDM. However, our analysis finds no significant statistical preference for any EDE model over $\Lambda$CDM. While some data combinations favour MSE over other EDE potentials (Tab. LABEL:tab:AIC), the improvement is marginal. Crucially, despite theoretical motivations for EDE models to reconcile the Hubble tension, observational data provide no compelling evidence supporting their necessity. Our constraints on $\log_{10}V_{0}$, $\mu$, $\phi_{0}$, and $f$ (Tab. LABEL:tab:fEDE) confirm that only a small EDE fraction is viable, except for the axionlike case. Moreover, the $H_{0}$ values computed using Eq. (36) align closely with direct constraints (Tab. LABEL:tab:constraint), validating Eq. (36) as a theoretical estimator. These results underscore the fundamental challenge of surpassing $\Lambda$CDM within the EDE framework. While EDE models allow for higher $H_{0}$ theoretically, the data do not support such an enhancement. This result is crucial for theoretical model building, as it suggests that conventional EDE scenarios, including the MSE potential, fail to resolve the Hubble tension.
In summary, we have introduced a unified model of EDE and LDE by modifying the SE (15) potential into the MSE potential (18), which allows for a CC like behaviour during recent times while exhibiting tracker dynamics during intermediate stages. We have compared this model with observational data alongside the SE, axionlike, and power law potentials. While all four potentials can theoretically produce an EDE solution that permits larger values of $H_{0}$, our analysis shows that, observationally, none of these models significantly improve the inferred value of $H_{0}$ compared to $\Lambda$CDM. Even in the case of the axionlike potential, where there is a slight increase in $H_{0}$, the improvement is not statistically significant. As a result, we find that EDE models, including the MSE potential, do not provide a compelling resolution to the Hubble tension, with $\Lambda$CDM remaining the most favoured model.
@@
Note Added: While working on the observational data analysis part of this paper we came across the recent paper [98] where the authors also have considered the MSE potential (18) but in the context of inflation.